Dr. Nick is twice as old as his son. Twenty years ago, he was 4 times as old as his son then. How old are they now?
Answer by mananth(16946)

let son's age be x
then father's age = 2x
20 years ago their ages
son (x-20)
father ---- (2x-20)
According to the information
(2x-20)=4(x-20)
2x-20 = 4x-80
2x=60
/2
x=30
Son's age = 30
Father's age = 30*2 = 60